# Transformer-based Wireless Symbol Detection Over Fading Channels

### v1

@@ VIII Conclusions
In this work, we proposed the DEFINED framework, inspired by decision feedback in wireless communication, to enhance symbol detection by incorporating decision pairs into prompts. Our approach achieved significant performance gains with limited pilot data while maintaining high accuracy with sufficient pilot data, making it highly adaptable for practical scenarios. Additionally, we introduced the IC-SSL model, demonstrating that Transformers could leverage both labeled data and unlabeled features for effective semi-supervised learning during inference. Extensive experiments across various modulation schemes in both SISO and MIMO systems validated the robustness and adaptability of our models. These contributions highlighted the potential of Transformers to perform both supervised and semi-supervised in-context learning, underscoring their suitability for future wireless communication systems. Furthermore, we conjecture that our DEFINED model learns the optimal semi-supervised estimator, a direction that warrants further theoretical investigation.

### v2

@@ VIII Conclusions
Inspired by decision feedback in wireless communication, we proposed DEFINED to enhance symbol detection by incorporating decision feedback into in-context prompts. Our approach achieved significant performance gains with limited pilot data while maintaining high accuracy when pilot information is sufficient, making it highly applicable to practical scenarios. Extensive experiments across various modulation schemes confirmed the robustness and adaptability of our model. Beyond empirical performance, we provided theoretical insights into the generalization behavior of pre-trained Transformers under SNR mismatch and derived an error lower bound for the DEFINED model. These findings highlight the potential of Transformer architectures as powerful in-context learners for wireless systems and motivate further theoretical exploration of their limits and capabilities.
Proof of Theorem 1.
We denote $\mu=\mu_{1}-\mu_{0}$, $u=2(\mu_{1}+\mu_{0})$, $q=y_{\mathsf{query}}$. Define $p=\frac{2}{k}\sum_{i=1}^{k}y_{i}x_{i}$. Since with probability $\mathbb{P}\left(x_{i}=1\right)=1/2$, $y_{i}=\mu_{1}+v_{i}$, with probability $\mathbb{P}\left(x_{i}=-1\right)=1/2$, $y_{i}=\mu_{0}+v_{i}$, where $v_{i}\sim\mathsf{N}(0,\Lambda)$, we have $p=2k_{1}\mu_{1}/k-2k_{0}\mu_{0}/k+g$, where $g=\frac{2}{k}\sum_{i=1}^{k}v_{i}$, $g\sim\mathsf{N}(0,4\Lambda/k)$, $k_{1}$ is a binomial random variable. Defining $h=k_{1}/k-1/2$, we have $p=\mu+hu+g$. The output of the trained Transformer is
$\widehat{x}_{\mathsf{out}}=S\left(\left(\frac{2}{k}\sum_{i=1}^{k}y_{i}x_{i}^{\top}\right)\Lambda^{-1}y_{\mathsf{query}}\right)=S(p^{\top}\Lambda^{-1}q).$ (9)
The probability of $x_{\mathsf{query}}=1$ given $y_{\mathsf{query}}$ is
$\mathbb{P}\left(x_{\mathsf{query}}=1|y_{\mathsf{query}}\right)=S((\mu_{1}-\mu_{0})^{\top}\Lambda^{-1}x_{\mathsf{query}})=S(\mu^{\top}\Lambda^{-1}q).$
Defining $a=\mu^{\top}\Lambda^{-1}q$, $b=(hu+g)^{\top}\Lambda^{-1}q$, we have
$p^{\top}\Lambda^{-1}q=(\mu+hu+g)^{\top}\Lambda^{-1}q$
$=$ $(hu+g)^{\top}\Lambda^{-1}q+\mu^{\top}\Lambda^{-1}q=a+b,$
and $S(p^{\top}\Lambda^{-1}q)=S(a+b)=S(a)+S^{\prime}(a)b+S^{\prime\prime}(\xi(a,b))b^{2}/2$, where $\xi$ are real numbers between $a$ and $a+b$. Thus, we have
$\mathbb{E}[\left(S(a+b)-S(a)\right)^{2}]=\mathbb{E}[\left(S^{\prime}(a)b+S^{\prime\prime}(\xi(a,b))b^{2}/2\right)^{2}]$
$=$ $\mathbb{E}[(S^{\prime}(a))^{2}b^{2}]+\mathbb{E}[S^{\prime}(a)S^{\prime\prime}(\xi(a,b))b^{3}+(S^{\prime\prime}(\xi(a,b)))^{2}b^{4}/4]$
Note that $\mathbb{E}[(S^{\prime}(a))^{2}b^{2}]=(S^{\prime}(a))^{2}\mathbb{E}[b^{2}]$. Let us first consider:
$\mathbb{E}[b^{2}]=$ $\mathbb{E}[h^{2}u^{\top}\Lambda^{-1}qu^{\top}\Lambda^{-1}q+g^{\top}\Lambda^{-1}qg^{\top}\Lambda^{-1}q]$
=$\overset{(a)}{=}$ $(u^{\top}\Lambda^{-1}q)^{2}/(4k)+4q^{\top}\Lambda^{-1}q/k$
where $(a)$ is due to Lemma D.1 in [41] that $\mathbb{E}[h^{2}]=1/(4k)$, $g^{\top}\Lambda^{-1}qg^{\top}\Lambda^{-1}q=q^{\top}\Lambda^{-1}gg^{\top}\Lambda^{-1}q=(gg^{\top}\Lambda^{-1}q)^{\top}\Lambda^{-1}q$ and $\mathbb{E}[gg^{\top}]=4\Lambda/k$.
For term $\mathbb{E}[S^{\prime}(a)S^{\prime\prime}(\xi(a,b))b^{3}+(S^{\prime\prime}(\xi(a,b)))^{2}b^{4}/4]$, we have
$\mathbb{E}[|S^{\prime}(a)S^{\prime\prime}(\xi(a,b))b^{3}+(S^{\prime\prime}(\xi(a,b)))^{2}b^{4}/4|]\leq$ $\mathbb{E}[|b^{3}|]+\mathbb{E}[b^{4}].$
Terms in $\mathbb{E}[|b^{3}|]$ contain three $h$, or three $g$, or one $h$ and two $g$, or two $h$ and one $g$. According to Lemma D.1, we have $\mathbb{E}[|h|]=O(1/\sqrt{k})$, $\mathbb{E}[h^{2}]=1/(4k)$, $\mathbb{E}[|h^{3}|]=O(k^{-3/2})$. Moreover, $g=2k^{-1/2}\Lambda^{1/2}\bar{g}$ and $\bar{g}\sim\mathsf{N}(0,I)$. Converting one $g$ to $\bar{g}$, we have a coefficient of $k^{-1/2}$. Thus, terms in $\mathbb{E}[|b^{3}|]$ are $O(k^{-3/2})$. Similarly, we have $\mathbb{E}[b^{4}]=O(k^{-2})$. Thus, we have
$\mathbb{E}[|S^{\prime}(a)S^{\prime\prime}(\xi(a,b))b^{3}+(S^{\prime\prime}(\xi(a,b)))^{2}b^{4}/4|]=o(1/k).$
Putting all together, we have:
$\mathbb{E}[|\widehat{x}_{\mathsf{out}}-\mathbb{P}\left(x_{\mathsf{query}}=1|y_{\mathsf{query}}\right)|^{2}]=\mathbb{E}[\left(S(a+b)-S(a)\right)^{2}]$
$=$ $\frac{1}{k}(S^{\prime}(a))^{2}[(u^{\top}\Lambda^{-1}q)^{2}/4+4q^{\top}\Lambda^{-1}q]+o(1/k).$
Proof of Theorem 2.
Suppose a random variable $x$ conditioned on $y$ has distribution $\mathbb{P}\left(x=1|y\right)=p$ and $\mathbb{P}\left(x=-1|y\right)=1-p$. Suppose the distribution of a predictor $\tilde{x}$ conditioned on $y$ is $\mathbb{P}\left(\tilde{x}=1|y\right)=q$ and $\mathbb{P}\left(\tilde{x}=-1|y\right)=1-q$. Given $y$, the accuracy between $x$ and $\tilde{x}$ is $Acc(x,\tilde{x})=pq+(1-p)(1-q)$, where $p,q\in[0,1]$. We can easily find that the best predictor $\tilde{x}$ that maximize the accuracy is
$\tilde{x}=\left\{\begin{array}[]{ll}1,&ifp>1/2\\ -1,&ifp\leq 1/2\end{array}\right.$
Note that in our setting, $\mathbb{P}\left(x_{\mathsf{query}}=1|y_{\mathsf{query}}\right)=S(\sigma^{-2}(\mu_{1}-\mu_{0})^{\top}y_{\mathsf{query}})$. Thus, the best predictor $x_{p}$ that maximize the accuracy between $x_{p}$ and $x_{\mathsf{query}}$ should be
$x_{p}=\left\{\begin{array}[]{ll}1,&if(\mu_{1}-\mu_{0})^{\top}y_{\mathsf{query}}>0\\ -1,&if(\mu_{1}-\mu_{0})^{\top}y_{\mathsf{query}}\leq 0\end{array}\right.$
Moreover, according to Theorem 1, when $k\to\infty$, $\widehat{x}_{\mathsf{out}}\to S(\xi^{-2}(\mu_{1}-\mu_{0})^{\top}y_{\mathsf{query}}).$ Thus, if we define $\widehat{x}_{p}$ as
$\widehat{x}_{p}=\left\{\begin{array}[]{ll}1,&ifx_{\mathsf{out}}>1/2\\ -1,&ifx_{\mathsf{out}}\leq 1/2\end{array}\right.$
Then, $\widehat{x}_{p}$ is equivalent to $x_{p}$ when $k\to\infty$. ∎